OpenStudy (handsomechickenlover):

What is the distance between the two stars? Show or explain your work. Star A: 0.176*10^18 Star D: 125.5*10^15

OpenStudy (handsomechickenlover):

Star A: \[0.176*10^{18}\]

OpenStudy (handsomechickenlover):

Star D: \[125.5*10^{15}\]

OpenStudy (handsomechickenlover):

HELP ME

OpenStudy (handsomechickenlover):

WILL MEDAL AND FAN

OpenStudy (handsomechickenlover):

...

OpenStudy (handsomechickenlover):

HALP ME PLEASEEEEE

OpenStudy (campbell_st):

well what you need to rewrite the numbers correctly in scientific notation \[0.176 \times 10^{18} = 1.76 \times 10^{17}\] scientific notation is a number between 1 and 10 that is then multiplied by a power of 10 looking at star D \[125.5 \times 10^{15} = 1.255 \times 10^{17}\] since to powers of 10 are the same, you can subtract, Any scientific calculator should easily do this problem.

OpenStudy (handsomechickenlover):

So, I subtract the 2 decimals?

OpenStudy (campbell_st):

no you are subtracting \[1.76 \times 10^{17} - 1.255 \times 10^{17} = ( 1.76 - 1.255)\times10^{17} \] then you will probably need to adjust the power of 10

OpenStudy (handsomechickenlover):

oh

OpenStudy (handsomechickenlover):

Thank you for your help :) heres your rewards!